OpenStudy (jiteshmeghwal9):

Here is a logical question of kindergarten. Please do not use calculator. The question is :- \(\Large{\frac{271}{9}=?}\)

OpenStudy (jiteshmeghwal9):

@satellite73 @ganeshie8 @AccessDenied @Chlorophyll @CliffSedge have a try :)

OpenStudy (jiteshmeghwal9):

@Zarkon @jim_thompson5910 @UnkleRhaukus @phi

jimthompson5910 (jim_thompson5910):

27/9 = 3 270/9 = 30 271/9 = 30 remainder 1

OpenStudy (anonymous):

Who needs a calculator? \(\large = \frac{270}{9}+\frac{1}{9}\)